What are the mole fractions of hydrochloric acid (HCI) and water in a 20% (w/w) aqeous HCl solution?

1
Expert's answer
2022-03-10T18:01:58-0500

mass of H2O =100gm-20gm=80gm

Mass of H2O=80gm

Now

"\frac{20gm\times1mol hcl}{36.46gm}=0.5485Hcl"

Now

"\frac{80gm\times1mol hcl}{18.015g}=4.4407molH_2o"

"\chi_{hcl}=\frac{mol hcl}{total mol}"

"\chi_{hcl}=\frac{0.5585}{0.548+4.4407}=0.1"


"\chi_{H_2o}=(1-\chi_{Hcl})=(1-0.1)=0.9"